# From Language to Action: A Review of Large Language Models as Autonomous Agents and Tool Users

### v1

@@ 8.2 Fine-tuning: embedding domain expertise and core behavioral traits
While prompt engineering allows for quick control during inference, fine-tuning offers an additional method to embed specialized knowledge and important traits into an agent’s design[39, 58, 36]. This process updates the model’s weights using curated datasets. It is essential for developing smaller, more efficient agents that can imitate the abilities of larger models or for teaching complex skills that are difficult to achieve through prompting alone[38]. Fine-tuning is the main way to adapt to new fields, gain skills, and improve the basic reasoning and interaction abilities of agents based on large language models[60, 74, 68]. A common practice in developing agents is using fine-tuning for knowledge distillation. In this method, the advanced reasoning of a large, skilled model generates high-quality training data for a smaller, open-source model[38]. This makes powerful capabilities more accessible. For example, Qin et al.[52] pioneered this with ToolLLaMA by fine-tuning LLaMA on a vast dataset of tool-use instructions and solutions from ChatGPT. This led to a highly capable open-source tool-using agent. Many have adopted this approach; Qian et al.[66] created Mistral-Interact by fine-tuning Mistral-7B on user dialogues simulated by GPT-4. Pang et al. applied the MATRIX framework to produce social simulation data for alignment fine-tuning. This method also works for specialized fields, as shown by Wu et al.[74] with ChatEDA, where a fine-tuned model, AutoMage, outperformed GPT-4 in electronic design automation tasks. The process is often iterative. Frameworks like Star-Agents from Zhou et al.[44] use multi-agent systems to automatically generate, evaluate, and refine instruction data to improve the fine-tuning process itself. Beyond simple knowledge transfer, fine-tuning is essential for developing the complex behavioral and cognitive skills that define an agent’s main abilities. This includes social intelligence, where Wang et al.[97] used data from guided social interactions to fine-tune SOTOPIA-$\pi$. This approach enabled a 7B model to achieve social goal completion similar to that of a larger GPT-4 agent. Fine-tuning can also improve self-awareness skills. Qu et al.[153] created RISE, an iterative fine-tuning process that helps agents to think about their own actions and fix their mistakes, which is difficult to accomplish through prompting alone. Similarly, Bo et al.[111] fine-tuned a shared "reflector" agent within their COPPER framework to enhance collaboration among multiple agents. This method works particularly well for embodied agents, as fine-tuning connects abstract reasoning with physical action. Zhai et al.[133] used reinforcement learning to fine-tune Vision-Language Models (VLMs), improving their decision-making in goal-oriented tasks. Yang et al.[123] employed an LLM expert to create distillation data to fine-tune a VLM called EMMA, teaching it to understand and interact with a visual environment. Fine-tuning is also crucial for transforming general models into specialized experts that can function in complex, real-world settings. In the realm of embodied AI, Schumann et al.[57] demonstrated that while prompting VELMA had some success, fine-tuning on navigation examples resulted in a 25% relative improvement in task completion. Hong et al.[59] fine-tuned CogAgent on GUI-grounding data, greatly enhancing its ability to navigate and manage computer interfaces. This specialization is vital in the fields of science and medicine. Liu et al.[79] fine-tuned DrBioRight 2.0 on a large cancer proteomics dataset to develop an expert bioinformatics chatbot. In addition, de Almeida et al.[136] fine-tuned ChatNT to create a multimodal agent with a deep understanding of DNA, RNA, and protein sequences. However, embedding expertise comes with risks; Yang et al.[93] showed that the fine-tuning process itself can be an avenue for backdoor attacks, where tainted data might train an agent to carry out harmful actions secretly. Ultimately, fine-tuning and prompting work together in a complementary way. Good prompting often serves as a foundation for creating the high-quality data needed for effective fine-tuning. On the other hand, fine-tuning can help agents respond better to complex prompts and improve their planning. Yin et al.[45] illustrate this with the LUMOS framework, demonstrating that fine-tuning on unified, high-quality annotations enhances both planning and grounding skills. This interaction facilitates the development of strong, specialized, and efficient agents where parametric training builds basic skills and non-parametric prompting guides their active use during inference.
@@ 8.3 Memory augmentation: enabling grounded reasoning and experiential learning
While prompting gives immediate instructions for an agent, memory augmentation allows agents to go beyond their temporary context limits[38, 37]. This enables grounded reasoning, long-term planning, and experiential learning[51, 68]. Unlike stateless agents that handle each turn independently, memory-augmented agents keep and use information over time. This changes them from simple instruction-followers into adaptable systems[71]. Memory can be divided into two main types: retrieval of external, fixed knowledge to ground reasoning in facts, and accumulation of internal, dynamic experience that supports learning and self-correction. The most common type of memory augmentation is Retrieval-Augmented Generation (RAG). This provides agents with access to external knowledge bases. RAG is crucial for grounding agent reasoning in specific, verifiable information. It helps reduce mistakes and improve reliability. For example, in clinical settings, frameworks like TRIAGEAGENT[38] and ColaCare[75] use RAG to give agents access to medical handbooks and guidelines. This ensures their decisions are based on solid evidence. This method works well in specialized technical areas; Barra[83] employs RAG to offer established simulation guidelines for scenario design. Xia et al.[72] use it to access technical datasheets for creating digital twin models. However, the effectiveness of RAG can depend on the model. Xia et al. mention a “cheat-sheet effect,” where it greatly improves performance in weaker LLMs compared to stronger ones. This external memory can also be dynamic, as shown in frameworks like DeCoAgent[37]. Here, a JSON memory module pulls current on-chain data, preventing unnecessary blockchain scans. Beyond retrieving static facts, memory is vital for experiential and reflective learning. It enables agents to improve through trial and error without needing updates to their parameters. One important framework in this area is Reflexion, proposed by Shinn et al.[13]. Here, an agent verbally reflects on task feedback and stores these reflections in an episodic memory buffer. This helps guide future attempts. This process of self-correction has shown to be effective in many applications, from improving public health policies in simulations[107] to generating high-quality distillation data for training other models[123]. The concept has evolved into more organized, long-lasting memory systems. Zhang et al.[68] introduce REMEMBERER, which updates an agent’s long-term experience memory through reinforcement learning. Wu et al.[51] implement a "Memory Bank" in AVATAR to store high-performing action sequences and instructions. Similarly, Star-Agents[44] uses an "Instruction Memory Bank" to continuously improve data generation strategies. Memory also helps agents stay grounded in their immediate, dynamic environment, especially in interactive settings. This is often done by updating the agent’s prompt with real-time perceptual information. For example, Song et al.[118] keep plans grounded in LLM-Planner by constantly updating the prompt with a list of visually perceived objects. Yang et al.[117] use prompts to coordinate visual grounding tools. This grounding includes more abstract representations of state as well. Li et al.[104] show that giving agents an explicit belief state representation greatly enhances planning and reduces errors in Theory of Mind tasks. In multi-agent systems, memory can be distributed. The CoA framework from Zhang et al.[41] passes a "Communication Unit" in sequence between agents, allowing each to build on the previous work. A more complex version is CoELA[60], which resembles human thinking with distinct semantic, episodic, and procedural memory modules to support long-term cooperation. Despite its advantages, memory enhancement is not perfect and faces significant challenges. The main limitation lies in how well the agent can manage long and complex memory streams. George et al.[101] points out that even with Chain-of-Thought prompting, agents can struggle to use crucial information from lengthy histories filled with distracting facts. Additionally, some memory strategies can backfire in specific situations. Zhang et al.[91] discovered that a Reflexion-style memory system led to worse performance in complex multitasking scenarios. Sometimes, simpler approaches like a "sliding window" memory can be a more effective, if less advanced, solution[90]. Ultimately, how well memory is implemented is what distinguishes an agent as a reactive tool versus a cognitive entity capable of learning, adapting, and maintaining coherent long-term behavior.
@@ 8.4 The synergistic integration of prompting, fine-tuning, and memory
The development of capable LLM-based agents relies on the complex interaction of three main mechanisms: dynamic, inference-time prompting; persistent, parametric fine-tuning; and context-aware memory augmentation[34, 35, 91]. These components work together, enabling the creation of agents that are both flexible and strong. Prompting offers a simple interface for real-time control and task specification. Fine-tuning provides permanent skills and domain-specific knowledge by adjusting model weights[38]. Memory, in its various forms, connects static knowledge with dynamic context, allowing agents to learn from experience and base their reasoning on relevant information. The combined use of these three elements is crucial for increasing agent autonomy, from single-agent reasoning to complex multi-agent collaboration[39, 34, 69]. At its core, prompt engineering is the main method for guiding agent behavior. It defines agent roles, organizes multi-agent systems, and allows tool use without changing the underlying model. In multi-agent setups, system-level prompts set the rules of engagement. For example, in MechAgents[34], roles like "planner" and "critic" are assigned through instructions, and in AGENTVERSE[35], expert roles are generated dynamically in a zero-shot manner. This approach works well in specialized areas; TRIAGEAGENT[38] and MAC[77] use role-playing prompts to simulate multidisciplinary medical teams, while frameworks like ChatSim[120] and ProtAgents[69] use structured prompts to manage collaborative workflows between agents that call external software tools. The accuracy of this control is vital. Target-driven planning prompts also enhance performance in complex environments. Similarly, frameworks like EASYTOOL[53] and ChemCrow[56] show that well-crafted prompts that clearly describe tool APIs are necessary for reliable tool use. However, relying solely on prompting can be fragile. Performance drops significantly when agents deal with long contexts filled with distracting "red herring" facts, even with CoT prompting. While prompting provides dynamic control, fine-tuning offers a method for permanently embedding specialized knowledge and complex skills in an agent. This is essential for creating agents that are not only trained by instructions but also inherently capable in specific domains. This is evident in the development of specialized models like DrBioRight 2.0[79] for bioinformatics, AutoMage[74] for electronic design automation, and CogAgent[59] for GUI navigation, all fine-tuned on domain-specific data. Besides domain knowledge, fine-tuning can build more abstract, general skills. For example, the RISE framework[153] uses an iterative fine-tuning process to teach agents the skill of self-correction. Meanwhile, SOTOPIA-$\pi$ [97] fine-tunes a model on social interaction data to enhance its social intelligence. The COPPER framework[111] fine-tunes a dedicated "reflector" agent to boost the collaborative abilities of an entire multi-agent system. This method is especially effective for grounding agents in new areas. For instance, VELMA[57] is fine-tuned to navigate street-view environments, and EMMA[123] fine-tunes a Vision-Language Model (VLM) using data from an LLM expert to help it interact with the visual world. One of the most effective combinations occurs when prompting serves as a tool for generating data, creating high-quality datasets for refining smaller, more efficient agents. This model lets the reasoning abilities of large proprietary models be distilled into open-source alternatives. A clear example is the ToolLLM framework [52], which uses ChatGPT to produce a vast dataset of instructions and API call solutions to fine-tune LLaMA into the highly effective ToolLLaMA. A similar method is used by Qian et al.[66] to develop Mistral-Interact by fine-tuning on user dialogues generated by prompting GPT-4. Pang et al.[98] also use the MATRIX framework to create alignment data. This method also applies to creating high-quality annotations for training. For instance, Yin et al.[45] employ few-shot prompts to format human annotations for the LUMOS framework, and Jorgensen[48] demonstrates it for generating entire training curricula for genetic programming agents. The Star-Agents framework[44] automates this whole process, using a multi-agent system to generate, evaluate, and improve instruction data for the best fine-tuning. This cycle, where a powerful teacher model generates data to refine a smaller student model, shows a scalable and effective way to make agent capabilities more accessible. Memory is essential for connecting an agent’s fixed skills with a changing environment, allowing for learning, growth, and grounded thinking. This enhancement has two main types: episodic, short-term memory for self-reflection and external, long-term memory for knowledge retrieval. The Reflexion framework[13] led the way in using verbal reinforcement, where an agent considers task feedback and saves this reflection in an episodic buffer to inform future attempts. This technique improves performance without needing weight updates. Behari[107] also uses this idea to adjust policy goals based on simulation results. However, both Zhang et al.[91] and Xing et al.[113] point out that reflection-based memory can hinder performance in complex multitasking or environmental navigation tasks. The second type, Retrieval-Augmented Generation (RAG), has become crucial for grounding agents in factual, domain-specific knowledge. By incorporating external knowledge bases, agents can bypass their internal knowledge limitations and lessen hallucinations. This is especially important in fields like medicine, where frameworks like TRIAGEAGENT[38], ColaCare[75], and the agent created by Ferber[76] use RAG to access medical handbooks and guidelines. This technique applies broadly, providing agents with on-chain data in decentralized systems[37], technical documents for AAS generation[72], and a knowledge base for personalized travel planning[67]. The sequential passing of a "Communication Unit" in the Chain-of-Agents framework[41] acts as a type of chained memory, allowing agents to collaborate and build on each other’s work in long-context tasks. The most advanced agents successfully combine all three components. In these systems, fine-tuning develops core skills, prompting guides inference-time reasoning and tool use, while memory offers dynamic context and knowledge. The CoELA framework[60] fine-tunes CoLLAMA using data collected by agents and applies structured prompts alongside a multi-part memory system (semantic, episodic, procedural) to enable complex collaboration. OmniJARVIS[62] demonstrates this integration by fine-tuning a VLA model on unified tokens that represent instructions, memories, and actions. These tokens are coordinated by prompts for open-world tasks. Similarly, Xu et al. suggest a split learning system that combines fine-tuned models on edge devices with RAG-augmented models on servers. This is managed through internal prompting. This three-part approach shows that while each part is effective on its own, bringing them together is essential for developing truly autonomous, capable, and reliable LLM-based agents.
@@ 9 Evaluation benchmarks and datasets
The rapid growth of LLM agents has required a change in how we evaluate them. Early methods that depended on static NLP benchmarks do not effectively capture the interactive, goal-driven, and often multi-step nature of these systems. The literature shows a clear shift toward more dynamic, specific, and analytical evaluation methods[91, 89, 97].
To give a clear overview of this progress, this section looks at the evaluation landscape through three connected lenses. First, it reviews the specialized benchmarks and interactive environments designed to test the abilities of agents in realistic situations[90, 38]. Next, it examines the various methods and metrics developed to judge performance beyond simple accuracy, including factors like efficiency, quality, and behavioral strength[90]. Finally, the section highlights the important role of the datasets used to train and ground these agents since they provide the base for agentic behavior.
@@ 9.1 Task-oriented and interactive benchmarks
Evaluating the capabilities of LLM-based agents requires shifting from static language metrics to dynamic benchmarks that focus on task performance and interactivity in complex environments. A notable trend is the creation of simulated environments where agents must engage in multi-step reasoning and actions. For example, TIME-ARENA[91] introduced a text-based simulation that includes time-based dynamics and constraints, challenging agents with multitasking scenarios in cooking, household tasks, and lab work. Advancing this complexity, AndroidArena[113] offers a general-purpose operating system environment to assess agents on intricate tasks that need cooperation between applications and compliance with user constraints. In a more specific area, RE-Bench provides demanding, open-ended machine learning research tasks, allowing for a direct comparison between AI agents and human experts. Environments like Overcooked-AI, ALFRED, and WebShop are established settings for testing zero-shot coordination, instruction following, and web navigation, respectively[25, 118, 36].
@@
Several meta-evaluation frameworks have been created to consolidate and analyze agent performance across various tasks. AgentBoard[90] offers an evaluation board that includes nine different multi-turn, partially observable tasks from areas like embodied AI, gaming, and web navigation. It provides a detailed metric to track progress, offering deeper insights beyond basic success rates. In the medical field, new benchmarks are emerging to ensure clinical safety and effectiveness. TRIAGE AGENT[38] released the first public benchmark for clinical triage, using metrics like discordance and undertriage rates against human expert performance. Other studies assess agents on their ability to execute clinical calculations using established medical calculators or make complex oncology decisions based on multimodal patient data[55, 76]. These interactive benchmarks are essential for fostering the growth of more capable, reliable, and aligned LLM agents.

### v2

@@ 8.2 Fine-tuning: embedding domain expertise and core behavioral traits
While prompt engineering allows for quick control during inference, fine-tuning offers an additional method to embed specialized knowledge and important traits into an agent’s design [38, 56, 35]. This process updates the model’s weights using curated datasets. It is essential for developing smaller, more efficient agents that can imitate the abilities of larger models or for teaching complex skills that are difficult to achieve through prompting alone [37]. Fine-tuning is the main way to adapt to new fields, gain skills, and improve the basic reasoning and interaction abilities of agents based on LLMs [58, 71, 65]. A common practice in developing agents is using fine-tuning for knowledge distillation. In this method, the advanced reasoning of a large, skilled model generates high-quality training data for a smaller, open-source model [37]. This makes powerful capabilities more accessible. For example, Qin et al. [50] pioneered this with ToolLLaMA by fine-tuning LLaMA on a vast dataset of tool-use instructions and solutions from ChatGPT. This led to a highly capable open-source tool-using agent. Qian et al. [63] created Mistral-Interact by fine-tuning Mistral-7B on user dialogues simulated by GPT-4. It focuses on enhancing user-agent interaction, proactively clarifying vague instructions, and refining user intentions before task execution. Pang et al. [94] applied the MATRIX framework to produce social simulation data for alignment fine-tuning. This method also works for specialized fields, as shown by Wu et al. [71] with ChatEDA, where a fine-tuned model, AutoMage, outperformed GPT-4 in electronic design automation tasks. The process is often iterative. Frameworks like Star-Agents from Zhou et al. [42] use multi-agent systems to automatically generate, evaluate, and refine instruction data to improve the fine-tuning process itself. Beyond simple knowledge transfer, fine-tuning is essential for developing the complex behavioral and cognitive skills that define an agent’s main abilities. This includes social intelligence, where Wang et al. [93] used data from guided social interactions to fine-tune SOTOPIA-$\pi$. This approach enabled a 7B model to achieve social goal completion similar to that of a larger GPT-4 agent. Fine-tuning can also improve self-awareness skills. Qu et al.[146] created RISE, an iterative fine-tuning process that helps agents to think about their own actions and fix their mistakes, which is difficult to accomplish through prompting alone. Similarly, Bo et al.[107] fine-tuned a shared "reflector" agent within their COPPER framework to enhance collaboration among multiple agents. This method works particularly well for embodied agents, as fine-tuning connects abstract reasoning with physical action. Zhai et al.[127] used RL to fine-tune VLMs, improving their decision-making in goal-oriented tasks. Yang et al.[119] employed an LLM expert to create distillation data to fine-tune a VLM called EMMA, teaching it to understand and interact with a visual environment. Fine-tuning is also crucial for transforming general models into specialized experts that can function in complex, real-world settings. In the realm of embodied AI, Schumann et al.[55] demonstrated that while prompting VELMA had some success, fine-tuning on navigation examples resulted in a 25% relative improvement in task completion. Hong et al.[57] fine-tuned CogAgent on GUI-grounding data, greatly enhancing its ability to navigate and manage computer interfaces. This specialization is vital in the fields of science and medicine. Liu et al.[76] fine-tuned DrBioRight 2.0 on a large cancer proteomics dataset to develop an expert bioinformatics chatbot. In addition, de Almeida et al.[130] fine-tuned ChatNT to create a multimodal agent with a deep understanding of DNA, RNA, and protein sequences. However, embedding expertise comes with risks; Yang et al.[89] showed that the fine-tuning process itself can be an avenue for backdoor attacks, where tainted data might train an agent to carry out harmful actions secretly. Ultimately, fine-tuning and prompting work together in a complementary way. Good prompting often serves as a foundation for creating the high-quality data needed for effective fine-tuning. On the other hand, fine-tuning can help agents respond better to complex prompts and improve their planning. Yin et al.[43] illustrate this with the LUMOS framework, demonstrating that fine-tuning on unified, high-quality annotations enhances both planning and grounding skills. This interaction facilitates the development of strong, specialized, and efficient agents where parametric training builds basic skills and non-parametric prompting guides their active use during inference.
Moreover, fine-tuning introduces risks, including model overfitting, amplification of biases present in the training data, and susceptibility to malicious data injection. While fine-tuning can achieve higher task-specific performance, it requires careful dataset curation, extensive computational resources, and rigorous validation. In contrast, prompt engineering offers flexibility and safety, but generally cannot reach the same depth of expertise.
@@ 8.3 Memory augmentation: enabling grounded reasoning and experiential learning
While prompting gives immediate instructions for an agent, memory augmentation extends agents beyond context limits[37, 36] by changing them from simple instruction-followers into adaptable systems[68]. Memory can be divided into two main types: retrieval of external knowledge (RAG) and accumulation of internal dynamic experience that supports learning and self-correction. RAG grounds agent reasoning in verifiable information. It helps reduce mistakes and improve reliability. For example, in clinical settings, frameworks like TRIAGEAGENT[37] and ColaCare[72] use RAG to give agents access to medical handbooks and guidelines. This ensures their decisions are based on solid evidence. This method works well in specialized technical areas; Barra[80] employs RAG to offer established simulation guidelines for scenario design. Xia et al.[69] use it to access technical datasheets for creating digital twin models. However, the effectiveness of RAG can depend on the model. Xia et al. mention a “cheat-sheet effect,” where it greatly improves performance in weaker LLMs compared to stronger ones. This suggests that the value of external memory depends on the agent’s inherent reasoning capabilities. The external memory can also be dynamic, as shown in frameworks like DeCoAgent[36]. Here, a JSON memory module pulls current on-chain data, preventing unnecessary blockchain scans. Beyond retrieving static facts, memory is vital for experiential and reflective learning. One important framework in this area is Reflexion[13] that stores verbal reflections in episodic buffers. This helps guide future attempts. This process of self-correction has been shown to be effective in many applications, from improving public health policies in simulations[103] to generating high-quality distillation data for training other models[119]. However, compared to an episodic system like Reflexion, RAG is more reliable for factual grounding but less effective for long-term adaptation. The concept has evolved into more organized, long-lasting memory systems. REMEMBERER[65] updates an agent’s long-term experience memory through RL. AVATAR [49] maintains a "Memory Bank" to store high-performing action sequences and instructions. Similarly, Star-Agents[42] uses an "Instruction Memory Bank" to continuously improve data generation strategies. Memory also helps agents stay grounded in their immediate, dynamic environment, especially in interactive settings. This is often done by updating the agent’s prompt with real-time perceptual information. For example, Song et al.[114] keep plans grounded in LLM-Planner by constantly updating the prompt with a list of visually perceived objects. Yang et al.[113] use prompts to coordinate visual grounding tools. This grounding includes more abstract representations of the state as well. Li et al.[100] show that giving agents an explicit belief state representation greatly enhances planning and reduces errors in Theory of Mind tasks. In multi-agent systems, memory can be distributed. The CoA framework from Zhang et al.[40] passes a "Communication Unit" in sequence between agents, allowing each to build on the previous work. A more complex version is CoELA[58], which resembles human thinking with distinct semantic, episodic, and procedural memory modules to support long-term cooperation. Despite its advantages, memory enhancement is not perfect and faces significant challenges. The main limitation lies in how well the agent can manage long and complex memory streams. George et al.[97] points out that even with CoT prompting, agents can struggle to use crucial information from lengthy histories filled with distracting facts. Additionally, some memory strategies can backfire in specific situations. Zhang et al.[87] discovered that a Reflexion-style memory system led to worse performance in complex multitasking scenarios. Sometimes, simpler approaches like a "sliding window" memory can be a more effective, if less advanced, solution[23]. Ultimately, how well memory is implemented is what distinguishes an agent as a reactive tool versus a cognitive entity capable of learning, adapting, and maintaining coherent long-term behavior.
@@ 8.4 The synergistic integration of prompting, fine-tuning, and memory
The development of capable LLM-based agents relies on the interaction of prompting, fine-tuning, and memory augmentation[33, 34, 87]. The combined use of these three elements is crucial for increasing agent autonomy, from single-agent reasoning to complex multi-agent collaboration[38, 33, 66]. At its core, prompt engineering is the main method for guiding agent behavior. However, relying solely on prompting can be fragile. Performance drops significantly when agents deal with long contexts filled with distracting "red herring" facts, even with CoT prompting. While prompting provides dynamic control, fine-tuning offers a method for permanently embedding specialized knowledge and complex skills in an agent. This is essential for creating agents that are trained by instructions and are inherently capable in specific domains. One of the most effective combinations occurs when prompting serves as a tool for generating data, creating high-quality datasets for refining smaller, more efficient agents. This model allows the reasoning abilities of large proprietary models to be distilled into open-source alternatives. A clear example is the ToolLLM framework [50], which uses ChatGPT to produce a vast dataset of instructions and API call solutions to fine-tune LLaMA into the highly effective ToolLLaMA. A similar method is used by Qian et al.[63] to develop Mistral-Interact by fine-tuning on user dialogues generated by prompting GPT-4. Pang et al.[94] also use the MATRIX framework to create alignment data. This bidirectional relationship extends further: fine-tuning can enhance an agent’s ability to follow complex prompts and execute domain-specific instructions more reliably. Yin et al.[43] demonstrate this with the LUMOS framework, showing that fine-tuning on unified, high-quality annotations improves both planning and grounding abilities when combined with structured prompting strategies.
Memory augmentation bridges prompting and fine-tuning mechanisms by providing the contextual foundation necessary for effective prompting and the experiential data required for targeted fine-tuning. RAG-based systems like TRIAGEAGENT[37] and ColaCare[72] use external memory to ground prompts in factual medical knowledge, mitigating hallucinations during inference. Conversely, episodic memory systems like Reflexion [13] accumulate task-specific experience in the form of episodic verbal reflections, which are used in subsequent prompts to improve both immediate reasoning and performance. The most advanced agents successfully combine all three components. In these systems, fine-tuning develops core skills, prompting guides inference-time reasoning and tool use, while memory offers dynamic context and knowledge. The CoELA framework[58] fine-tunes CoLLAMA using data collected by agents and applies structured prompts alongside a multi-part memory system (semantic, episodic, procedural) to enable complex collaboration. OmniJARVIS[60] demonstrates this integration by fine-tuning a VLA model on unified tokens that represent instructions, memories, and actions. These tokens are coordinated by prompts for open-world tasks. These systems illustrate that, while each mechanism is effective independently, their combination is essential for developing truly autonomous, capable, and reliable LLM-based agents.
@@ 9 Evaluation benchmarks and datasets
The rapid growth of LLM agents has required a change in how we evaluate them. Early methods that depended on static NLP benchmarks do not effectively capture the interactive, goal-driven, and often multi-step nature of these systems. The literature shows a clear shift towards more dynamic, specific, and analytical evaluation methods[87, 86, 93].
To give a clear overview of this progress, this section looks at the evaluation landscape through three connected lenses. First, it reviews the specialized benchmarks and interactive environments designed to test the abilities of agents in realistic situations[23, 37]. Next, it examines the various methods and metrics developed to judge performance beyond simple accuracy, including factors like efficiency, quality, and behavioral strength[23]. Finally, the section highlights the important role of the datasets used to train and ground these agents since they provide the basis for agentic behavior.
@@ 9.1 Task-oriented and interactive benchmarks
Evaluating the capabilities of LLM-based agents requires a shift from static language metrics to dynamic benchmarks that focus on task performance and interactivity in complex environments. A notable trend is the creation of simulated environments where agents must engage in multi-step reasoning and actions. For example, TIME-ARENA[87] introduced a text-based simulation that includes time-based dynamics and constraints, challenging agents with multitasking scenarios in cooking, household tasks, and lab work. Advancing this complexity, AndroidArena[109] offers a general-purpose operating system environment to assess agents on intricate tasks that need cooperation between applications and compliance with user constraints. In a more specific area, RE-Bench provides demanding, open-ended machine learning research tasks, allowing for a direct comparison between AI agents and human experts. Environments like Overcooked-AI, ALFRED, and WebShop are established settings for testing zero-shot coordination, instruction following, and web navigation, respectively [25, 114, 35].
@@
Several meta-evaluation frameworks have been created to consolidate and analyze agent performance across various tasks. AgentBoard[23] offers an evaluation board that includes nine different multi-turn, partially observable tasks from areas like embodied AI, gaming, and web navigation. It provides a detailed metric to track progress, offering deeper insights beyond basic success rates. In the medical field, new benchmarks are emerging to ensure clinical safety and effectiveness. TRIAGE AGENT[37] released the first public benchmark for clinical triage, using metrics like discordance and undertriage rates against human expert performance. Other studies assess agents in their ability to execute clinical calculations using established medical calculators or make complex oncology decisions based on multimodal patient data[53, 73]. These interactive benchmarks are essential for fostering the growth of more capable, reliable, and aligned LLM agents.